The squared visibility amplitude is estimated using

V^{2} = \frac{< X^{2} + Y^{2} − N >}{< N >^{2}},

where X and Y are the real and imaginary parts of the complex visibility (derived by Fourier transform of the bin counts at the fringe frequency), and N is the total count rate (sum of the bin count rates). The brackets denote average. We have left off a leading constant, which is related to the reduction in amplitude due to the binning.